OpenStudy (anonymous):

how to find the circumference of a circle if the radius is 2x+1?

OpenStudy (anonymous):

plug it into the magical formula, of course!

OpenStudy (anonymous):

C=2*3.14*r 6.28*2x +1....so now what :( am i even close?

jimthompson5910 (jim_thompson5910):

it helps to place 'r' in parenthesis \[\Large C = 2\pi r\] \[\Large C = 2\pi ( r)\] \[\Large C = 2\pi (2x+1)\]

jimthompson5910 (jim_thompson5910):

when you say 6.28*2x +1, you're only multiplying the 6.28 by the 2x when it should be 6.28 times (2x+1)

OpenStudy (anonymous):

awesome!! thank you!!!